# On Phase Unwrapping via Digital Wavefront Sensors

### v2

@@ 3.3 The digital Shack-Hartmann wavefront sensor
$E^{m}_{j,k}(\boldsymbol{x},\phi):=\mathcal{F}^{-1}(\chi_{0,0}(\boldsymbol{\xi})\mathcal{F}(E^{m}(\cdot,\phi))(T_{j,k}(\boldsymbol{\xi})))(\boldsymbol{x})\,,\qquad\forall\boldsymbol{x}\in\mathbb{R}^{2}\,.$ (3.6)
@@
This assertion follows directly from [19, Proposition 3.8] using $\mathcal{F}(E)\equiv 1$. ∎
From (3.8) one can see that the intensity images on each subaperture of the CCD detector of the SH-WFS are essentially shifted versions of each other, and that the relative shifts $(\boldsymbol{s}_{j,k}-\boldsymbol{s}_{0,0})$ correspond to the average slopes of the local wavefront aberration. These relative shifts can e.g. be estimated via matching the subimages $\left|E^{m}_{j,k}\right|^{2}$ and $\left|E^{m}_{0,0}\right|$; see e.g. [26, 27, 19]. Alternatively, as suggested in [19] they can also be calculated via
@@ 3.4 Digital Fourier-type wavefront sensors
$\operatorname{OTF}_{\psi}(\boldsymbol{\xi}):=e^{i\psi(\boldsymbol{\xi})}\,.$
@@
$\psi(\boldsymbol{\xi}_{1},\boldsymbol{\xi}_{2})=$ $c(|\boldsymbol{\xi}_{1}|+|\boldsymbol{\xi}_{2}|)$ $c|\boldsymbol{\xi}_{1}|$$\begin{cases}-2c\boldsymbol{\xi}_{1}\,,&\frac{-\pi}{3}<\arctan{\frac{\boldsymbol{\xi}_{1}}{\boldsymbol{\xi}_{2}}}<\frac{\pi}{3}\\ c(\boldsymbol{\xi}_{1}-\sqrt{3}\boldsymbol{\xi}_{2})\,,&\frac{\pi}{3}<\arctan{\frac{\boldsymbol{\xi}_{1}}{\boldsymbol{\xi}_{2}}}<\pi\\ c(\boldsymbol{\xi}_{1}+\sqrt{3}\boldsymbol{\xi}_{2})\,,&\text{otherwise}\end{cases}$
@@
$\psi(\boldsymbol{\xi}_{1},\boldsymbol{\xi}_{2})=$ $c\sqrt{|\boldsymbol{\xi}_{1}|^{2}+|\boldsymbol{\xi}_{2}|^{2}}$ $c|\boldsymbol{\xi}_{2}|$$\begin{cases}\frac{\pi}{2}\,,&\boldsymbol{\xi}_{1}\boldsymbol{\xi}_{2}<0\,,\\ 0\,,&\text{otherwise}\end{cases}$
Table 3.1: Shape functions $\psi$ corresponding to the Fourier-type WFSs depicted in Figure 2.2. The constant $c>0$ relates to the apex angle of the pyramidal prisms. The typical choice $c\approx 1$ is used in the numerical examples presented in Section 4.
@@
$E(\boldsymbol{x})=E(\boldsymbol{x},\phi):=\chi_{\Omega}(\boldsymbol{x})e^{-i\phi(\boldsymbol{x})}\,,$
where $\phi$ are the wavefront aberrations to be reconstructed, and the domain $\Omega$ corresponds to the aperture of the imaging system. This is physically realized if an additional lens is placed in front of the imaging system in Figure 2.2, and if the incoming light stems from a sufficiently distant point source. In this case, (3.10) becomes
$E^{m}(\boldsymbol{x})=E^{m}_{\psi}(\boldsymbol{x},\phi)=\mathcal{F}^{-1}\left(e^{i\psi(\boldsymbol{\xi})}\mathcal{F}\left(\chi_{\Omega}e^{-i\phi}\right)(\boldsymbol{\xi})\right)(\boldsymbol{x})\,.$ (3.11)
Finally, the camera/detector in the conjugate plane of the imaging system in Figure 2.2 measures the intensity of the field $E^{m}$, i.e.,
@@
For the reconstruction of the aberration $\phi$ from the intensity measurements $I_{\psi}$, a number of different reconstruction algorithms have been proposed. Most of these depend heavily on the shape of the chosen optical element, i.e., on the concrete form of $\psi$, and are often based on linearizations of (3.12). An overview of reconstruction methods for the most commonly used 4-sided PWFS can e.g. be found in [44]. Additionally, AI-based methods such as Deep Optics PWFS have recently been proposed [51]. For arbitrary choices of $\psi$, including those which do not correspond to a pyramidal prism but can e.g. be realized by a spatial light modulator, the NOnlinear Pyramid Extension (NOPE) [21, 20] has been proposed.
@@
Among the Fourier-type WFSs, the 4-sided PWFS is the most commonly used. In [45], the Preprocessed Cumulative Reconstructor with Domain Decomposition (PCuReD), based on CuReD for the SH-WFS discussed above, is suggested as a suitable and highly efficient reconstructor, which gives rise to the following algorithm.
@@
    $I^{m}_{\psi}(\boldsymbol{x},\phi_{w})=\left|\mathcal{F}^{-1}\left(e^{ic(|\boldsymbol{\xi}_{1}|+|\boldsymbol{\xi}_{2}|)}\mathcal{F}\left(\chi_{\Omega}e^{-i\phi_{w}}\right)(\boldsymbol{\xi})\right)(\boldsymbol{x})\right|^{2}\,.$
@@ 4 Numerical experiments
In FSOC, laser beams are used for the wireless transmission of data in e.g. optical laser communications between satellites and ground-based stations. Thanks to high bandwidths, laser communication provides large and fast data transfer rates. However, when the laser signal travels through the Earth’s atmosphere, it gets disrupted by turbulence. Recent research in FSOC, see e.g. [31, 8, 47, 38, 50], focuses on using AO to overcome this problem. In the design phase of AO systems for FSOC, simulations of laser light propagation including diffraction and interference are required. To this end, physical optics propagation is commonly employed [3], which uses Fourier transforms to compute phase angles. As these phase angles contain $2\pi$ ambiguities, fast and accurate phase unwrapping algorithms are of great interest for solving this problem.
4.1 Implementation and algorithms
@@
The digital WFS phase unwrapping algorithms described in Section 3 were all implemented in MATLAB. As mentioned above, for the digital SH-WFS based unwrapping the CuReD method is used in Step 4 of Algorithm 2. This algorithm requires neither parameter tuning nor an initial guess [57, 43, 42]. Similarly, for Step 2 in Algorithm 4 the PCuReD algorithm is used. Finally, in Step 2 of Algorithm 3 the NOPE algorithm is used, which does require an initial guess, for which either zero, or an approximation computed via PCuReD has been used [21]. This is denoted below by “zero starting” and “linear starting”, respectively. Furthermore, the NOPE also has an optional smoothness parameter $s$, which, unless noted otherwise, is set to the canonical choice $s=11/6$ as suggested in [21]; see also [25]. In order to validate the performance of our proposed algorithms, we compare them to a number of commonly used local and global phase unwrapping approaches. Among the local methods we use both the built-in MATLAB unwrapping function, as well as the more advanced algorithm proposed in [18]. The latter is based on unwrapping most reliable pixels (MRP), and has a MATLAB implementation available in [11]. Moreover, we use two global methods based on solving the Poisson equation [55] and the transport of intensity (TIE) equation [56], respectively. The corresponding MATLAB codes are available in [58] and [59].
@@ 4.2 Numerical results
For the first set of numerical tests, we consider a random phase $\phi$ created by the well-known OCTOPUS simulation tool of the European Southern Observatory [28], corresponding to the wavefront aberration caused by some atmospheric turbulence profile. Figure 4.1 contains both the actual phase $\phi$, the manually wrapped phase $\phi_{w}$, and its noisy version $\phi_{w}^{\delta}$ containing $20\%$ relative random noise. The noise was added after the wrapping, which accounts for the larger range of values of $\phi_{w}^{\delta}$ in comparison with $\phi_{w}$. This noisy wrapped phase $\phi_{w}^{\delta}$ was then used as input for the unwrapping algorithms.
@@ 30 $\times$ 30 100 $\times$ 100 500 $\times$ 500
Finally, we compare the results of our proposed digital WFS unwrapping approaches with those obtained by four different state-of-the art methods introduced above. The corresponding results are depicted in Figure 4.5. First, note that the built-in Matlab unwrapping function fails entirely, while the MRP method produces discontinuous, patch-like reconstructions which are also far off from the true phase. The best reconstruction is the one obtained via the PE method, which is similar to, but less detailed than that of the digital SH-WFS unwrapping. The TIE unwrapping captures many of the structural features of the ground truth phase, but some features almost appear to be inverted with respect to the ground truth. In summary, we find that our proposed digital SH-WFS and Fourier-type WFS methods perform at least as good as, and often much better than, these four standard approaches. In terms of reconstruction quality, the digital SH-WFS with a suitable number of subapertures, and the digital 4-sided PWFS using the NOPE with linear starting produce the best results among our methods.
@@
For the second example, we consider a phase unwrapping problem from FSOC, in particular the Advanced Laser Guide Star Adaptive Optics for Satellite Communication Assessments (ALASCA) project [1] currently under development, for which we now provide some brief background. Within the ALASCA project, an optical communication channel from the optical ground station of the European Space Agency (ESA) in Tenerife to communication satellites is developed. The concept is based on laser guide star AO and aims for 24/7 operations. In the design phase of the project, reliable simulations of the laser guide star AO system are crucial. Hence, the laser propagation in a turbulent atmosphere has to be modelled, which usually involves physical optics propagation methods and thus phase screens that contain $2\pi$ ambiguities. The simulations for ALASCA are carried out via the PyrAmid Simulator Software for Adaptive opTics Arcetri (PASSATA)[2], which is IDL and CUDA based object oriented software for Monte-Carlo end-to-end AO simulations. Figure 4.6 shows a typical wrapped phase screen occurring during daytime conditions, i.e., strong turbulence, simulated with PASSATA, for which no ground truth unwrapped phase is available.
@@ 6 Support
This research was funded in part by the Austrian Science Fund (FWF) SFB 10.55776/F68 “Tomography Across the Scales”, project F6805-N36 (Tomography in Astronomy) and project F6807-N36 (Tomography with Uncertainties). For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. Furthermore, the authors were supported by the Austrian Research Promotion Agency (FFG) project number FO999888133, as well as the NVIDIA Corporation Academic Hardware Grant Program.

### v4

@@ 3.3 The digital Shack-Hartmann wavefront sensor
$E^{m}_{j,k}(\boldsymbol{x},\phi):=\mathcal{F}^{-1}\left(\chi_{0,0}(\boldsymbol{\xi})\mathcal{F}(E^{m}(\cdot,\phi))(T_{j,k}(\boldsymbol{\xi}))\right)(\boldsymbol{x})\,,\qquad\forall\boldsymbol{x}\in\mathbb{R}^{2}\,.$ (3.6)
@@
This assertion follows directly from [22, Proposition 3.8] using $\mathcal{F}(E)\equiv 1$. ∎
From (3.8) one can see that the intensity images on each subaperture of the CCD detector of the SH-WFS are essentially shifted versions of each other, and that the relative shifts $(\boldsymbol{s}_{j,k}-\boldsymbol{s}_{0,0})$ correspond to the average slopes of the local wavefront aberration. These relative shifts can e.g. be estimated via matching the subimages $\left|E^{m}_{j,k}\right|$ and $\left|E^{m}_{0,0}\right|$; see e.g. [28, 29, 22]. Alternatively, as suggested in [22] they can also be calculated via
@@ 3.4 Digital Fourier-type wavefront sensors
$\operatorname{OTF}_{\psi}(\boldsymbol{\xi}):=e^{i\psi(\boldsymbol{\xi})}\,.$ (3.11)
@@
$\psi(\xi_{1},\xi_{2})=$ $c(|\xi_{1}|+|\xi_{2}|)$ $c|\xi_{1}|$$\begin{cases}-2c\xi_{1}\,,&\frac{-\pi}{3}<\arctan{\frac{\xi_{1}}{\xi_{2}}}<\frac{\pi}{3}\\ c(\xi_{1}-\sqrt{3}\xi_{2})\,,&\frac{\pi}{3}<\arctan{\frac{\xi_{1}}{\xi_{2}}}<\pi\\ c(\xi_{1}+\sqrt{3}\xi_{2})\,,&\text{otherwise}\end{cases}$
@@
$\psi(\xi_{1},\xi_{2})=$ $c\sqrt{|\xi_{1}|^{2}+|\xi_{2}|^{2}}$ $c|\xi_{2}|$$\begin{cases}\frac{\pi}{2}\,,&\xi_{1}\xi_{2}<0\,,\\ 0\,,&\text{otherwise}\end{cases}$
Table 3.1: Shape functions $\psi(\boldsymbol{\xi})$, with $\boldsymbol{\xi}=(\xi_{1},\xi_{2})$, corresponding to the Fourier-type WFSs depicted in Figure 2.2. The constant $c>0$ relates to the apex angle of the pyramidal prisms. The typical choice $c\approx 1$ is used in the numerical examples presented in Section 4.
@@
$E(\boldsymbol{x})=E(\boldsymbol{x},\phi):=\chi_{\Omega}(\boldsymbol{x})e^{-i\phi(\boldsymbol{x})}\,,$ (3.12)
where $\phi$ is the wavefront aberration to be reconstructed, and the domain $\Omega$ corresponds to the aperture of the imaging system. This is physically realized by an additional lens placed in front of the imaging system in Figure 2.2 in case the incoming light stems from a sufficiently distant point source. With this, (3.10) becomes
$E^{m}_{\psi}(\boldsymbol{x},\phi):=E^{m}(\boldsymbol{x})=\mathcal{F}^{-1}\left(e^{i\psi(\boldsymbol{\xi})}\mathcal{F}\left(\chi_{\Omega}e^{-i\phi}\right)(\boldsymbol{\xi})\right)(\boldsymbol{x})\,,$ (3.13)
where we have used (3.11) and (3.12). Finally, the camera/detector in the conjugate plane of the imaging system in Figure 2.2 measures the intensity of the field $E^{m}$, i.e.,
@@
For the reconstruction of the aberration $\phi$ from the intensity measurements $I_{\psi}$, a number of different reconstruction algorithms have been proposed. Most of these depend heavily on the shape of the chosen optical element, i.e., on the concrete form of $\psi$, and are often based on linearizations of (3.14). An overview of reconstruction methods for the most commonly used 4-sided PWFS can, e.g., be found in [46]. Additionally, AI-based methods such as Deep Optics PWFS have recently been proposed [53]. For arbitrary choices of $\psi$, including those which do not correspond to a pyramidal prism but can, e.g., be realized by a spatial light modulator, the NOnlinear Pyramid Extension (NOPE) [23] has been proposed.
@@
As noted above, the 4-sided PWFS is the most commonly used Fourier-type WFS in practice. In [47], the Preprocessed Cumulative Reconstructor with Domain Decomposition (PCuReD), based on CuReD for the SH-WFS discussed above, is suggested as a suitable and highly efficient reconstructor, which gives rise to the following algorithm.
@@
Given the wrapped phase $\phi_{w}(\boldsymbol{\xi})$ for all $\boldsymbol{\xi}=(\xi_{1},\xi_{2})\in\Omega\subset{\mathbb{R}^{2}}$, compute the unwrapped phase $\phi_{r}(\boldsymbol{\xi})$ via:
@@
    $I^{m}_{\psi}(\boldsymbol{x},\phi_{w})=\left|\mathcal{F}^{-1}\left(e^{ic(|\xi_{1}|+|\xi_{2}|)}\mathcal{F}\left(\chi_{\Omega}e^{-i\phi_{w}}\right)(\boldsymbol{\xi})\right)(\boldsymbol{x})\right|^{2}\,.$
@@ 4 Numerical experiments
In FSOC, laser beams are used for the wireless transmission of data in, e.g., optical laser communications between satellites and ground-based stations. Thanks to high bandwidths, laser communication provides large and fast data transfer rates. However, when the laser signal travels through the Earth’s atmosphere, it gets disrupted by turbulence. Recent research in FSOC focuses on using AO to overcome this problem; see e.g. [33, 10, 49, 41, 52]. In the design phase of AO systems for FSOC, simulations of laser light propagation including diffraction and interference are required. To this end, physical optics propagation is commonly employed [3], which uses Fourier transforms to compute phase angles. As these phase angles contain $2\pi$ ambiguities, fast and accurate phase unwrapping algorithms are of great interest for solving this problem.
@@
The digital WFS phase unwrapping algorithms described in Section 3 were all implemented in MATLAB. As mentioned above, for the digital SH-WFS based unwrapping the CuReD method is used in Step 4 of Algorithm 2. This algorithm requires neither parameter tuning nor an initial guess [65, 45, 44]. Similarly, for Step 2 in Algorithm 4 the PCuReD algorithm is used. Finally, in Step 2 of Algorithm 3, the NOPE algorithm is used, which does require an initial guess, for which either zero, or an approximation computed via PCuReD has been used [23]. This is denoted below by “zero starting” and “linear starting”, respectively. Furthermore, the NOPE also has an optional smoothness parameter $s$, which, unless noted otherwise, is set to the canonical choice $s=11/6$ as suggested in [23]; see also [27].
4.1 Comparison with existing methods
To validate the performance of our proposed algorithms, we compare them to a number of commonly used local and global phase unwrapping approaches, which are based on line integration of an estimate of the phase gradient. While the former evaluate the integral along a single path, the latter average integrals along different paths.
Among the local methods, we use the built-in MATLAB unwrapping function, as well as the algorithm proposed in [20]. The MATLAB built-in unwrapping works in one-dimension. If the input is a matrix, the function operates column-wise. The phase is unwrapped by adding multiples of $2\pi$ to the phase whenever the jump between two consecutive pixels is greater than or equal to $\pi$ radians. The method proposed in [20] is more advanced, and is based on unwrapping most reliable pixels (MRP). Its MATLAB implementation is available in [13]. As a first step, the algorithm calculates the reliability of each pixel using second order derivatives. Then, it follows a non-continuous path to perform the unwrapping process, in which pixels with higher reliability are unwrapped first. This algorithm is robust, fast, and is used inside AO simulation tools. However, in situations where the noise level is high the quality of the unwrapping suffers.
In contrast to local methods, global methods utilize the wrapped values together with some additional (global) assumptions, such as the probability distribution of the noise or the regularity of the phase field. Among those methods, we select one algorithm based on the PE [63] and one based on the TIE [64]. For both methods, MATLAB codes are available in [61] and [62], respectively. In these approaches, first an auxiliary complex field is generated using the wrapped phase. Then, a two-dimensional second-order elliptic partial differential equation (PDE) is solved, which uses the longitudinal intensity derivative of this field as an input. As the names suggest, in the case of the PE algorithm the Poisson equation is solved, while in the case of the TIE the transport of intensity equation is solved. The resulting phase profile is then automatically in unwrapped form, since it has been obtained as the solution of a PDE rather than as the argument of a complex-valued function. Those approaches have already been successfully applied in, e.g., interferometry and digital holography applications [34].
In recent years, deep learning methods have become more and more popular for phase unwrapping. Promising results have been shown in the literature; see e.g. [60, 55, 21]. However, the training phase of neuronal networks requires a lot of data samples, experience, and time. Moreover, it is not so straight forward to include the neuronal network within the AO simulation tool we are using, and would also require a new network training process adapted to our discussed application. Hence, we do not further consider those approaches in the framework of this paper.
@@ 4.2 Numerical results
For the first set of numerical tests, we consider a random phase $\phi$ created by the OCTOPUS simulation tool (popular in the AO community) of the European Southern Observatory [30], corresponding to the wavefront aberration caused by some atmospheric turbulence profile. Figure 4.1 contains both the actual phase $\phi$, the manually wrapped phase $\phi_{w}$ created using the MATLAB built-in function wrapToPi(), and its noisy version $\phi_{w}^{\delta}$ containing $20\%$ relative uniform random noise. The noise was added after the wrapping, which accounts for the larger range of values of $\phi_{w}^{\delta}$ in comparison with $\phi_{w}$. This noisy wrapped phase $\phi_{w}^{\delta}$ was then used as input for the unwrapping algorithms.
@@ 30 $\times$ 30 100 $\times$ 100 500 $\times$ 500
Finally, we compare the results of our proposed digital WFS unwrapping approaches with those obtained by four different state-of-the art methods introduced above. The corresponding results are depicted in Figure 4.5. First, note that the built-in Matlab unwrapping function works only into one dimension. Depending on the given data, it may be sufficient to perform unwrapping in the two dimensions consecutively. However, in our case this entirely failed. The MRP method produces discontinuous, patch-like reconstructions which are also far off from the true phase. The best reconstruction is the one obtained via the PE method, which is similar to, but less detailed than that of the digital SH-WFS unwrapping. The TIE unwrapping captures many of the structural features of the ground truth phase, but some features almost appear to be inverted with respect to the ground truth. In summary, we find that our proposed digital SH-WFS and Fourier-type WFS methods perform at least as good as, and often much better than, these four standard approaches. Table 4.1 quantifies the reconstruction quality in terms of the relative error, structural similarity index (SSIM) [56] and multiscale SSIM (MS-SSIM) [57] for each of the considered approaches. Overall, we find that the digital SH-WFS with a suitable number of subapertures, and the digital 4-sided PWFS using the NOPE algorithm with linear starting produce the best results among all methods.
Digital Shack-Hartmann WFS Digital 4-sided Pyramid WFS
Criterion $30\times 30$ $100\times 100$ $500\times 500$ PCuReD NOPE NOPE+Lin. NOPE+Lin.($s=0.5$)
Rel.Error 41.51% 41.38% 193.87% 27.92% 67.37% 11.19% 30.8%
SSIM 0.5208 0.6119 0.4103 0.5554 0.5224 0.4527 0.5236
MS-SSIM 0.6576 0.8154 0.3654 0.7109 0.7168 0.7800 0.7120
Figure 4.2 Figure 4.3
Digital Fourier-type WFS Other Algorithms
Criterion 3-sided Pyramid Roof Cone iQuad MATLAB MRP PE TIE
Rel.Error 73.27% 86.18% 89.96% 98.75% 724.96% 417.43% 66.94% 81.17%
SSIM 0.5119 0.4629 0.4233 0.4121 0.4080 0.4171 0.4438 0.4431
MS-SSIM 0.7018 0.6439 0.4583 0.4671 0.3771 0.5101 0.6376 0.5885
Figure 4.4 Figure 4.5
Table 4.1: Comparison of the relative error, structural similarity index (SSIM) [56] and multiscale SSIM (MS-SSIM) [57] for each of the considered unwrapping approaches.
@@
For the second example, we consider a phase unwrapping problem of FSOC, in particular from the Advanced Laser Guide Star Adaptive Optics for Satellite Communication Assessments (ALASCA) project [1] currently under development, for which we now provide some brief background. Within the ALASCA project, an optical communication channel from the optical ground station of the European Space Agency (ESA) in Tenerife to communication satellites is developed. The concept is based on laser guide star AO and aims for 24/7 operations. In the design phase of the project, reliable simulations of the laser guide star AO system are crucial. Hence, the laser propagation in a turbulent atmosphere has to be modelled. This usually involves physical optics propagation methods and thus phase screens which contain $2\pi$ ambiguities. The simulations for ALASCA are carried out via the PyrAmid Simulator Software for Adaptive opTics Arcetri (PASSATA)[2], which is IDL and CUDA based object oriented software for Monte-Carlo end-to-end AO simulations. Figure 4.6 shows a typical wrapped phase screen occurring during daytime conditions, i.e., strong turbulence, simulated with PASSATA, for which no ground truth unwrapped phase is available.
@@ 6 Support
This research was funded in part by the Austrian Science Fund (FWF) SFB 10.55776/F68 “Tomography Across the Scales”, project F6805-N36 (Tomography in Astronomy) and project F6807-N36 (Tomography with Uncertainties), as well as 10.55776/P34981 “New Inverse Problems of Super-Resolved Microscopy (NIPSUM)”. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. Furthermore, the authors were supported by the Austrian Research Promotion Agency (FFG) project number FO999888133, as well as the NVIDIA Corporation Academic Hardware Grant Program. The financial support by the Austrian Federal Ministry for Digital and Economic Affairs, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged.